# On the Martin boundary for discrete TASEP

### v1

@@ 1.1. Overview
When $p=0$, our Gibbs probability measures are in bijection with coherent systems on (positive part of) the Gelfand–Tsetlin graph. In turn, the latter are in correspondence with characters of the infinite-dimensional unitary group $U(\infty)$, its spherical and finite-factor representations, as well as with infinite totally-positive Toeplitz matrices, see [Ols16, Section 9], [Gor21, Section 20.3]. Because of these connections, the $p=0$ classification problem is very well-studied, see [Edr53, Voi76, VK82, Boy83, OO98, BO12, Pet14, GP15] for various approaches. The $p=0$ problem can also be reformulated as a study of all possible limits of normalized Schur polynomials $s_{\lambda}$ as the number of variables grows to infinity:
@@
Two deformations of (1.1) were explored in the literature: the first one depends on a real parameter $\theta>0$ and replaces Schur polynomials with Jack polynomials, motivated by connections to spherical representations of Gelfand pairs at $\theta=1/2,1,2$ and to log-gases and $\beta$-ensembles of the random matrix theory. The answer in the Jack-deformed problem turns out to be very similar to the Schur $\theta=1$ case, see [OO98]. Another qq–deformation replaces $1$s in (1.1) with geometric series with denominator $q>0$ and relates to representation theory of quantum groups [Sat19, Sat21] and to $q^{\text{volume}}$–weighted random plane partitions [Gor12]. Here the role of $q$ turned out to be more significant and the answer for $q=1$ is very different from the general $q$ case of [Gor12, GO16]. The two deformations were subsequently lifted to a common generalization related to principal specializations of Macdonald polynomials in [Cue18, Ols21].
From the asymptotic representation theory perspective, in this paper we initiate the study of another deformation of (1.1), related to Grothendieck polynomials $G_{\lambda}$, see e.g. [FK94, FK96, IN13, MS13, Yel16, HJK+24] among many papers on these interesting polynomials. In contrast to all the previously studied cases from the last paragraph, the denominator in our version of (1.1) no longer has an explicit fully factorized form, which suggests that less formulas are available for our pp–deformation and leads us to develop alternative methods.
Switching to the statistical mechanics point of view, we recall that configurations of the celebrated six-vertex model (see [LW72, Bax07, Res10, BL14, GN23] for the reviews) assign to each vertex of a domain $\Omega\subset\mathbb{Z}^{2}$ one of the six types shown in Figure 1, in a way that is globally consistent: the result must form non-intersecting, possibly touching paths that connect specified boundary points of $\Omega$. We assign a positive weight $w_{i}>0$ to each vertex type $i=1,\dots,6$, typically denoted by $(a_{1},a_{2},b_{1},b_{2},c_{1},c_{2})$, and say that a probability measure on configurations in $\Omega$ is Gibbs, if for any finite subdomain $\Omega^{\prime}\subset\Omega$ and any boundary conditions — configuration of boundary points $\partial\Omega^{\prime}$ which the paths should connect, the conditional distribution of configurations $\sigma$ inside $\Omega^{\prime}$ has the form
@@
where $N_{i}(\sigma)$ is the number of type $i$ vertices in configuration $\sigma$ and $Z$ is a normalizing constant. The formula (1.2) silently assumes that the particular boundary condition on $\partial\Omega^{\prime}$ has a non-zero probability. We remark that the Gibbs property (1.2) depends on two, rather than six parameters due to four conservation laws, see [GN23, Lemma 2.1].
@@
We conjecture that the answer should be very different depending on the value of $\Delta:=\frac{a_{1}a_{2}+b_{1}b_{2}-c_{1}c_{2}}{2\sqrt{a_{1}a_{2}b_{1}b_{2}}}$. Namely, for $\Delta<1$, we expect the only measure to be the unit mass on the configurations of all paths going straight to the right. On the other hand, for $\Delta>1$, we expect a rich family of Gibbs measures to exist. The prediction is based on a phase transition at $\Delta=1$ discovered in [GL23].
In this paper, we deal with a particular case of Question 6v corresponding to $\Delta=+\infty$. Namely, we prohibit the $b_{2}$ vertex degenerating into the five vertex model, and arrange the weights of the remaining vertices as in the right panel of Figure 1. This is an $\varepsilon\to 0+$ limit of the stochastic weights (cf. [GS92, BCG16]) $(1,1,p,\varepsilon,1-p,1-\varepsilon)$ with $\Delta=\frac{p+\varepsilon}{2\sqrt{p\varepsilon(1-p)(1-\varepsilon)}}$. The Gibbs property for this instance of the five-vertex model can be also interpreted through transitional probabilities of discrete time TASEP with geometric jumps, see Section 2.5 for further details. Various questions about such TASEP were previously investigated e.g. in [DW08, MR23, MS13, KPS19]. In contrast to the general six-vertex model, our analysis in this situation is simplified by the existence of determinantal formulas for transition probabilities and connection to the Grothendieck polynomials.
In line with conjectures about Question 6v, our main result is the construction of a rich class of Gibbs measures in the quadrant for the five vertex model, see Theorem 3.14 for further details. Some of these measures have probabilistic interpretations as time evolutions of TASEP started from the step initial condition and with prescribed asymptotic speeds of individual particles, see Theorem 5.1. The phenomenon of asymptotic particle speeds is also known in $p=0$ case of the Gelfand-Tsetlin graph, since [VK82], which interpreted the parameters of Gibbs measures as normalized asymptotic lengths (i.e. speeds of growth) of rows and columns of corresponding randomly growing Young diagrams. However, there is a striking difference: in the Gelfand-Tsetlin graph any positive speeds are possible, but in our pp–deformation we discovered only speeds larger than $\frac{p}{1-p}$ to be admissible.
@@
Classifications of Gibbs measures and coherent systems have two important consequences in the context of statistical mechanics and integrable probability. First, knowing Gibbs measures in specific geometries leads to predictions for the local limits of models in various domains. Along these lines, [She05] classified all translationally invariance Gibbs measures on lozenge tilings in terms of their slopes and subsequently [Agg23] (see also [Gor17]) proved that these measures are the only possible bulk limits for uniformly random tilings of arbitrary domains. [OV96] classified Gibbs measures on spectra of corners of random matrices; based on that [OR06] predicted the universality of GUE-corners process as a limit of statistical mechanics models near boundaries; [AG22] proved this for lozenge tilings (and universality is expected to extend beyond, e.g., to the six-vertex model, cf. [GN23, GL23]). Okounkov and Sheffield predicted and [CH14, AH23] identified the Airy2 line ensemble as a particular solution to a classification problem involving Brownian Gibbs property for a family of continuous curves, which was subsequently used as a tool for proving convergence towards the Airy2 line ensemble. Similarly, we expect that our Gibbs measures (and, looking further ahead, the eventual answer to Question 6v) will describe possible local limits in the six-vertex model and its five-vertex degeneration.
Second, the Gibbs measures appearing as answers in the classification problems in infinite domains often turn out to be exactly solvable: many more formulas and algebraic structures are available for them, as compared to generic models of 2d statistical mechanics. For instance, the measures for the $p=0$ of our problem corresponding to the Gelfand-Tsetlin graph enjoy connections to determinantal point processes (e.g.  [BK08]), Robinson-Schensted-Knuth correspondence (e.g. [BBB+18] and references therein), to 2+1–dimensional interacting particle systems (e.g. [BF14]). Similarly, we expect that the answers to Question 6v and its five-vertex version are very special and worth further studies.
@@ 2. Grothendieck functions
In this section we describe the necessary facts about (stable) Grothendieck symmetric functions. We are using the version of these functions from [Yel16], which is our primary reference, together with [HJK+24]. For this section we let $\mathfrak{a},\mathfrak{b}$ denote a pair of generic parameters.
@@ 2.2. Grothendieck functions
We can also define $g_{\lambda}^{(\mathfrak{a},\mathfrak{b})}$ in a fashion similar to (2.1), see [HJK+24, Definition 1.2]. We only give the $\mathfrak{b}=0$ case of that definition.
Proposition 2.1 ([HJK+24]).
@@ 2.3. Branching, Jacobi-Trudi and skew Cauchy identities
For both $G^{(-\mathfrak{a},-\mathfrak{b})}_{\lambda/\mu}$ and $g^{(-\mathfrak{a},-\mathfrak{b})}_{\lambda/\mu}$ there are various analogues of Jacobi-Trudi identities, see [HJK+24]. In this work we need only two particular expressions described below. For both statements we use $f(\mathfrak{a}^{n})$ to denote the evaluation of $f$ at $n$-copies of $\mathfrak{a}$.
Proposition 2.6 ([HJK+24, Theorem 5.2]).
@@ 3.1. Definition of the graph
Straightforward modification of [Ols03, Proposition 10.3]. ∎
@@
Then $\{M^{\omega}_{n}\}_{n}$ is an extreme coherent system on $\mathbb{G}^{0}$ and $\Omega$ and every extreme coherent system is described in this way. See [OO98, BO12, Pet14, GP15].
@@ 3.2. Coherent systems $M^{\Phi}_{n}$
Our coherent systems are parametrized by complex functions from a space $\mathcal{F}$. This space consists of complex functions $\Phi(z)$ such that there exists a sequence $\Phi_{k}(z)$ satisfying
  • $\Phi_{k}(z)$ converge to $\Phi(z)$ uniformly on a neighborhood of the unit disk;
@@
Note that $\Phi(z)\in\mathcal{F}$ is analytic on the unit disk (the poles of $\Phi_{k}$ are at $\frac{1+y_{i,k}}{y_{i,k}}>1$) and $\Phi(1)=1$.
For $n>0$, $\lambda\in\mathbb{G}^{p}_{n}$ we define $M_{n}^{\Phi}(\lambda)$ by the following identity:
@@
Also set $M^{\Phi}_{0}(\varnothing):=1$. We postpone the convergence questions until later. For now, we treat the identity above as an identity of formal power series in $z_{1},\dots,z_{n}$, where $\Phi(z)\in\mathcal{F}$ is identified with its Taylor expansion at 0. We also should explain why $M_{n}^{\Phi}(\lambda)$ are well-defined, which is done in Proposition 3.8 below. More generally, for $\lambda,\mu\in\mathbb{G}_{n}^{p}$ we define $M_{n}^{\Phi}(\lambda/\mu)$ by
@@
Fix $\mu\in\mathbb{G}^{p}_{n}$. Note that $\Phi(z_{1})\Phi(z_{2})\dots\Phi(z_{n})G_{\mu}^{(0,-p)}(z_{1},\dots,z_{n})$ is a symmetric power series in $z_{1},\dots,z_{n}$. Since $\{s_{\lambda}(z_{1},\dots,z_{n})\}_{\lambda\in\mathbb{G}^{p}_{n}}$ form a basis of symmetric polynomials in $z_{1},\dots,z_{n}$, we have a decomposition of the form
$\Phi(z_{1})\Phi(z_{2})\dots\Phi(z_{n})G_{\mu}^{(0,-p)}(z_{1},\dots,z_{n})=\sum_{\lambda\in\mathbb{G}^{p}_{n}}c_{\lambda}s_{\lambda}(z_{1},\dots,z_{n}).$
@@
$M_{n}^{\Phi}(\lambda/\mu)=G_{\lambda}^{(0,-p)}(1^{n})\sum_{\nu:|\nu|\leq|\lambda|}c_{\nu}a_{\nu\lambda}.$
For the last statement note that the smallest degree of $\Phi(z_{1})\dots\Phi(z_{n})G_{\mu}^{(0,-p)}(z_{1},\dots,z_{n})$ is $|\mu|$, so $c_{\lambda}=0$ unless $|\lambda|\geq|\mu|$. Thus, the sum above computing $M_{n}^{\Phi}(\lambda/\mu)$ has only vanishing terms when $|\lambda|<|\mu|$. ∎
@@
For the next property assume that $\Phi(0)\neq 0$. Define an algebra homomorphism $\iota_{\Phi}:\Lambda\to\mathbb{C}$ by defining it on the generators $h_{k}$ in the following way:
@@
where the sum converges uniformly for $z$ in a neighborhood of the unit disk.
@@
From (2.3) we have $|G^{(0,-p)}_{\lambda/\mu}(z)|\leq z^{|\lambda|-|\mu|}$ when $|z|<p^{-1}$, so convergence reduces to the convergence of
@@
This last series converges uniformly and absolutely on a neighborhood of the unit disk since this is a Taylor expansion of $\Phi(z)$. So, both sides of (3.5) are analytic functions on the unit disk with the same Taylor expansions at 0 due to the formal identities. ∎

### v2

@@ 1.1. Overview
When $p=0$, our Gibbs probability measures are in bijection with coherent systems on (positive part of) the Gelfand–Tsetlin graph. In turn, the latter are in correspondence with characters of the infinite-dimensional unitary group $U(\infty)$, its spherical and finite-factor representations, as well as with infinite totally-positive Toeplitz matrices, see [olshanski2016markov, Section 9], [Gorin_book, Section 20.3]. Because of these connections, the $p=0$ classification problem is very well-studied, see [edrei1953generating, voiculescu1976representations, vershik1982characters, boyer1983infinite, okounkov1998asymptotics, borodin2012boundary, petrov2014boundary, gorin2015asymptotics] for various approaches. The $p=0$ problem can also be reformulated as a study of all possible limits of normalized Schur polynomials $s_{\lambda}$ as the number of variables grows to infinity:
@@
Two deformations of (1.1) were explored in the literature: the first one depends on a real parameter $\theta>0$ and replaces Schur polynomials with Jack polynomials, motivated by connections to spherical representations of Gelfand pairs at $\theta=1/2,1,2$ and to log-gases and $\beta$-ensembles of the random matrix theory. The answer in the Jack-deformed problem turns out to be very similar to the Schur $\theta=1$ case, see [okounkov1998asymptotics]. Another qq–deformation replaces $1$s in (1.1) with geometric series with denominator $q>0$ and relates to representation theory of quantum groups [sato2019quantized, sato2021inductive] and to $q^{\text{volume}}$–weighted random plane partitions [gorin2012q]. Here the role of $q$ turned out to be more significant and the answer for $q=1$ is very different from the general $q$ case of [gorin2012q, gorin2016quantization]. The two deformations were subsequently lifted to a common generalization related to principal specializations of Macdonald polynomials in [cuenca2018asymptotic, olshanski2021macdonald].
From the asymptotic representation theory perspective, in this paper we initiate the study of another deformation of (1.1), related to Grothendieck polynomials $G_{\lambda}$, see e.g. [fomin1994grothendieck, fomin1996yang, ikeda2013k, motegi2013vertex, Yel16, HJKSS21] among many papers on these interesting polynomials. In contrast to all the previously studied cases from the last paragraph, the denominator in our version of (1.1) no longer has an explicit fully factorized form, which suggests that less formulas are available for our pp–deformation and leads us to develop alternative methods.
Switching to the statistical mechanics point of view, we recall that configurations of the celebrated six-vertex model (see [Lieb_ferroelectric_models, baxter2007exactly, reshetikhin2010lectures, Bleher-Liechty14, Gorin_Nicoletti_lectures] for the reviews) assign to each vertex of a domain $\Omega\subset\mathbb{Z}^{2}$ one of the six types shown in Figure 1, in a way that is globally consistent: the result must form non-intersecting, possibly touching paths that connect specified boundary points of $\Omega$. We assign a positive weight $w_{i}>0$ to each vertex type $i=1,\dots,6$, typically denoted by $(a_{1},a_{2},b_{1},b_{2},c_{1},c_{2})$, and say that a probability measure on configurations in $\Omega$ is Gibbs, if for any finite subdomain $\Omega^{\prime}\subset\Omega$ and any boundary conditions — configuration of boundary points $\partial\Omega^{\prime}$ which the paths should connect, the conditional distribution of configurations $\sigma$ inside $\Omega^{\prime}$ has the form
@@
where $N_{i}(\sigma)$ is the number of type $i$ vertices in configuration $\sigma$ and $Z$ is a normalizing constant. The formula (1.2) silently assumes that the particular boundary condition on $\partial\Omega^{\prime}$ has a non-zero probability. We remark that the Gibbs property (1.2) depends on two, rather than six parameters due to four conservation laws, see [Gorin_Nicoletti_lectures, Lemma 2.1].
@@
We conjecture that the answer should be very different depending on the value of $\Delta:=\frac{a_{1}a_{2}+b_{1}b_{2}-c_{1}c_{2}}{2\sqrt{a_{1}a_{2}b_{1}b_{2}}}$. Namely, for $\Delta<1$, we expect the only measure to be the unit mass on the configurations of all paths going straight to the right. On the other hand, for $\Delta>1$, we expect a rich family of Gibbs measures to exist. The prediction is based on a phase transition at $\Delta=1$ discovered in [Gorin_Liechty_2023].
In this paper, we deal with a particular case of Question 6v corresponding to $\Delta=+\infty$. Namely, we prohibit the $b_{2}$ vertex degenerating into the five vertex model, and arrange the weights of the remaining vertices as in the right panel of Figure 1. This is an $\varepsilon\to 0+$ limit of the stochastic weights (cf. [gwa1992six, borodin2016stochastic]) $(1,1,p,\varepsilon,1-p,1-\varepsilon)$ with $\Delta=\frac{p+\varepsilon}{2\sqrt{p\varepsilon(1-p)(1-\varepsilon)}}$. The Gibbs property for this instance of the five-vertex model can be also interpreted through transitional probabilities of discrete time TASEP with geometric jumps, see Section 2.5 for further details. Various questions about such TASEP were previously investigated e.g. in [dieker2008determinantal, matetski2023tasep, motegi2013vertex, knizel2019generalizations]. In contrast to the general six-vertex model, our analysis in this situation is simplified by the existence of determinantal formulas for transition probabilities and connection to the Grothendieck polynomials.
In line with conjectures about Question 6v, our main result is the construction of a rich class of Gibbs measures in the quadrant for the five vertex model, see Theorem 3.15 for further details. Some of these measures have probabilistic interpretations as time evolutions of TASEP started from the step initial condition and with prescribed asymptotic speeds of individual particles, see Theorem 5.1. The phenomenon of asymptotic particle speeds is also known in $p=0$ case of the Gelfand-Tsetlin graph, since [vershik1982characters], which interpreted the parameters of Gibbs measures as normalized asymptotic lengths (i.e. speeds of growth) of rows and columns of corresponding randomly growing Young diagrams. However, there is a striking difference: in the Gelfand-Tsetlin graph any positive speeds are possible, but in our pp–deformation we discovered only speeds larger than $\frac{p}{1-p}$ to be admissible.
@@
Classifications of Gibbs measures and coherent systems have two important consequences in the context of statistical mechanics and integrable probability. First, knowing Gibbs measures in specific geometries leads to predictions for the local limits of models in various domains. Along these lines, [sheffield2005random] classified all translationally invariance Gibbs measures on lozenge tilings in terms of their slopes and subsequently [aggarwal2023universality] (see also [gorin2017bulk]) proved that these measures are the only possible bulk limits for uniformly random tilings of arbitrary domains. [olshanski1996ergodic] classified Gibbs measures on spectra of corners of random matrices; based on that [okounkov2006birth] predicted the universality of GUE-corners process as a limit of statistical mechanics models near boundaries; [aggarwal2022gaussian] proved this for lozenge tilings (and universality is expected to extend beyond, e.g., to the six-vertex model, cf. [Gorin_Nicoletti_lectures, Gorin_Liechty_2023]). Okounkov and Sheffield predicted and [corwin2014brownian, aggarwal2023strong] identified the Airy2 line ensemble as a particular solution to a classification problem involving Brownian Gibbs property for a family of continuous curves, which was subsequently used as a tool for proving convergence towards the Airy2 line ensemble. Similarly, we expect that our Gibbs measures (and, looking further ahead, the eventual answer to Question 6v) will describe possible local limits in the six-vertex model and its five-vertex degeneration.
Second, the Gibbs measures appearing as answers in the classification problems in infinite domains often turn out to be exactly solvable: many more formulas and algebraic structures are available for them, as compared to generic models of 2d statistical mechanics. For instance, the measures for the $p=0$ of our problem corresponding to the Gelfand-Tsetlin graph enjoy connections to determinantal point processes (e.g.  [borodin2008asymptotics]), Robinson-Schensted-Knuth correspondence (e.g. [betea2018perfect] and references therein), to 2+1–dimensional interacting particle systems (e.g. [borodin2014anisotropic]). Similarly, we expect that the answers to Question 6v and its five-vertex version are very special and worth further studies.
@@ 2. Grothendieck functions
In this section we describe the necessary facts about (stable) Grothendieck symmetric functions. We are using the version of these functions from [Yel16], which is our primary reference, together with [HJKSS21]. For this section we let $\mathfrak{a},\mathfrak{b}$ denote a pair of generic parameters.
@@ 2.2. Grothendieck functions
We can also define $g_{\lambda}^{(\mathfrak{a},\mathfrak{b})}$ in a fashion similar to (2.1), see [HJKSS21, Definition 1.2]. We only give the $\mathfrak{b}=0$ case of that definition.
Proposition 2.1 ([HJKSS21]).
@@ 2.3. Branching, Jacobi-Trudi and skew Cauchy identities
For both $G^{(-\mathfrak{a},-\mathfrak{b})}_{\lambda/\mu}$ and $g^{(-\mathfrak{a},-\mathfrak{b})}_{\lambda/\mu}$ there are various analogues of Jacobi-Trudi identities, see [HJKSS21]. In this work we need only two particular expressions described below. For both statements we use $f(\mathfrak{a}^{n})$ to denote the evaluation of $f$ at $n$-copies of $\mathfrak{a}$.
Proposition 2.6 ([HJKSS21, Theorem 5.2]).
@@ 3.1. Definition of the graph
Straightforward modification of [Olsh01, Proposition 10.3]. ∎
@@
Then $\{M^{\omega}_{n}\}_{n}$ is an extreme coherent system on $\mathbb{G}^{0}$ and $\Omega$ and every extreme coherent system is described in this way. See [okounkov1998asymptotics, borodin2012boundary, petrov2014boundary, gorin2015asymptotics].
@@ 3.2. Coherent systems $M^{\Phi}_{n}$
Let $\overline{\mathbb{D}}$ denote the closed unit disk $\{z\in\mathbb{C}:|z|\leq 1\}$. Let $\mathcal{F}$ be the space of complex functions $\Phi(z)$ on $\overline{\mathbb{D}}$ such that there exists a sequence $\Phi_{k}(z)$ satisfying
  • $\Phi_{k}(z)$ are complex functions converging to $\Phi(z)$ uniformly on $\overline{\mathbb{D}}$;
@@
Note that $\Phi(z)\in\mathcal{F}$ is analytic on the unit disk (the poles of $\Phi_{k}$ are at $\frac{1+y_{i,k}}{y_{i,k}}>1$) and $\Phi(1)=1$. We equip $\mathcal{F}$ with the topology of uniform convergence on $\overline{\mathbb{D}}$, which makes $\mathcal{F}$ a complete metric space.
For $n>0$, $\lambda\in\mathbb{G}^{p}_{n},\Phi\in\mathcal{F}$ we define $M_{n}^{\Phi}(\lambda)$ by the following identity:
@@
Also set $M^{\Phi}_{0}(\varnothing):=1$. We postpone the convergence questions until later. For now, we treat the identity above as an identity of formal power series in $z_{1},\dots,z_{n}$, where $\Phi(z)\in\mathcal{F}$ is identified with its Taylor expansion at 0. We also should explain why $M_{n}^{\Phi}(\lambda)$ are well-defined, which is done in Proposition 3.9. More generally, for $\lambda,\mu\in\mathbb{G}_{n}^{p}$ we define $M_{n}^{\Phi}(\lambda/\mu)$ by
@@
Let $\Phi\in\mathcal{F}$. Then the Taylor expansion of $\Phi$ at 0 has non-negative coefficients and converges uniformly on $\overline{\mathbb{D}}$.
Proof.
First note that for $x\in[-1,\frac{p}{1-p}]$, $y\geq 0$, $y\geq x$ and $z\in\overline{\mathbb{D}}$ we have
$\frac{1-x(z-1)}{1-y(z-1)}=\frac{1+x}{1+y}+\sum_{n\geq 1}\frac{(y-x)y^{n-1}}{(1+y)^{n+1}}z^{n},$
where all the coefficients $\frac{(y-x)y^{n-1}}{(1+y)^{n+1}}$ are non-negative. Hence for the function $\Phi_{k}(z)$ from (3.3) the Taylor expansion at 0 is of the form
$\Phi_{k}(z)=\sum_{n\geq 0}a_{n,k}z^{n},$
where $a_{n,k}\geq 0$ and the series converges for $|z|\leq 1$. Since $\Phi_{k}(1)=1$ we also know that $\sum_{n\geq 0}a_{n,k}=1$.
Let $\Phi\in\mathcal{F}$. Then there is a sequence $(\Phi_{k})_{k}$ of the form (3.3) converging uniformly to $\Phi$. By Cauchy’s integral formula the Taylor expansions
$\Phi_{k}(z)=\sum_{n\geq 0}a_{n}z^{n},\qquad\Phi_{k}(z)=\sum_{n\geq 0}a_{n,k}z^{n}$
satisfy $a_{n}=\lim_{k\to\infty}a_{n,k}\geq 0$ and $\sum_{n\geq 0}a_{n}\leq\lim_{k\to\infty}\sum_{n\geq 0}a_{n,k}=1$. This implies that $\sum_{n\geq 0}a_{n}z^{n}$ converges uniformly on $\overline{\mathbb{D}}$. ∎
Proposition 3.9.
@@
Fix $\mu\in\mathbb{G}^{p}_{n}$. Note that $\Phi(z_{1})\Phi(z_{2})\dots\Phi(z_{n})G_{\mu}^{(0,-p)}(z_{1},\dots,z_{n})$ is a symmetric power series in $z_{1},\dots,z_{n}$. Since the Schur polynomials $\{s_{\lambda}(z_{1},\dots,z_{n})\}_{\lambda\in\mathbb{G}^{p}_{n}}$ form a basis of symmetric polynomials in $z_{1},\dots,z_{n}$, we have a decomposition of the form
$\Phi(z_{1})\Phi(z_{2})\dots\Phi(z_{n})G_{\mu}^{(0,-p)}(z_{1},\dots,z_{n})=\sum_{\lambda\in\mathbb{G}^{p}_{n}}c^{\Phi}_{\lambda\mu}s_{\lambda}(z_{1},\dots,z_{n}).$
@@
$M_{n}^{\Phi}(\lambda/\mu)=G_{\lambda}^{(0,-p)}(1^{n})\sum_{\nu:|\nu|\leq|\lambda|}c^{\Phi}_{\nu\mu}a_{\nu\lambda}.$
For the last statement note that the smallest degree of $\Phi(z_{1})\dots\Phi(z_{n})G_{\mu}^{(0,-p)}(z_{1},\dots,z_{n})$ is $|\mu|$, so $c^{\Phi}_{\lambda,\mu}=0$ unless $|\lambda|\geq|\mu|$. Thus, in the sum above computing $M_{n}^{\Phi}(\lambda/\mu)$ all terms vanish when $|\lambda|<|\mu|$. ∎
@@
For the next property assume that $\Phi(0)\neq 0$. Define an algebra homomorphism $\iota_{\Phi}:\Lambda\to\mathbb{C}$ by its values on the generators $h_{k}$:
@@
where the sum converges uniformly on $\overline{\mathbb{D}}$.
@@
From (2.3) we have $|G^{(0,-p)}_{\lambda/\mu}(z)|\leq z^{|\lambda|-|\mu|}$ when $|z|\leq 1<p^{-1}$, so the convergence of (3.6) reduces to the convergence of
@@
This last series converges uniformly and absolutely on $\overline{\mathbb{D}}$ by Proposition 3.8. So, both sides of (3.6) are analytic functions on the unit disk with the same Taylor expansions at 0 due to the formal identities. ∎